OpenStudy (openst):

inverse tangent sqrt3/3

OpenStudy (openst):

\[\tan^{-1} \frac{ \sqrt{3} }{ 3 }\] How?

OpenStudy (prathamesh_m):

\[\frac{ \sqrt{3} }{ 3 } = \frac{ 1 }{ \sqrt{3} }\]

OpenStudy (prathamesh_m):

Can you find the answer now?

OpenStudy (phi):

***How?*** in general, you use a calculator. However, there are some angles that people memorize the sine , cosine and tangent they are 0º, 30º, 45º, 60º and 90º see http://www.analyzemath.com/trigonometry/special_angles.html in your problem, you have one of these "special" angles